[Brahms-l] Shift report 20050221 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Mon Feb 21 2005 - 16:03:43 EST
Shifter leader was CN
Shifters were 


1. good beam. Took data for about 6 hours
2.runs
MRS@40A1050, FS@8B2442 -
runno.(tr2, tr3) : 13351(85k, 149k), 13352(105k, 504k), 13353(73k,511k), 13356(89k, 153k) 
-
non-TPC pedestal runs (13348, 13354)
TPC pedestal runs (133349, 13355)
3. no detector problems.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Feb 21 16:05:16 2005

This archive was generated by hypermail 2.1.8 : Mon Feb 21 2005 - 16:05:28 EST